Using the cut-and-try method for aggregate operations planning, we can calculate theending inventory and then calculate the safety stock as a percent of forecast demand. Suppose that the beginning inventory is 300, the production requirement in units of product is 1,350, demand forecast is 1,500, what is the ending inventory and percent safety stock?

Multiple Choice

200 and 10%

150 and 10%

300 and 20%

450 and 30%

150 and 20%
Answer #1

Beginning inventory = 300

Production requirement = 1350

Demand forecast = 1500

Safety stock = production requirement - Demand forecast + beginning inventory

Let, safety stock = x%

So, 1500*(x/100) = 1350-1500+300

or, 15x = 150

or, x = 10

So, safety stock = 10%

Ending inventory = production requirement + beginning inventory - demand forecast = 1350+300-1500 = 150

So, the answer is 150 and 10%(Ans)